OpenStudy (bbb911):

This is Geometry: Suppose ^ABC ~ ^GHI with scale factor 3:5, AB = 9, BC = 18, and AC = 21. Find the perimeter of ^ GHI.

OpenStudy (anonymous):

pri of ABC=9+18+21=48 so GHI= 48*5/3=80

OpenStudy (bbb911):

:D